# Hyperbolic Network Latent Space Model with Learnable Curvature

### v1

@@ 4 Theoretical Analysis
In this section, we first utilize hyperbolic geometry techniques to characterize the error bound for the embedded distance matrix of latent positions from hyperbolic spaces with different curvatures, which reveals the importance of the latent space curvature for general distance-based hyperbolic embedding methods. We then present the consistency rates for the maximum likelihood estimators of both the curvature parameter and the latent positions.
We use $P$, $Z$, and similar symbols, to denote the true values of parameters, and use $\hat{P}$, $\hat{Z}$, and so on, to represent their estimators. The following assumptions are made throughout the analysis.
@@
    Values of both the estimators and true model parameters belong to the bounded parameter space
    $\Xi=\left\{0<C_{1}^{-1}\leq K\leq C_{1},0<C_{2}\leq|\Theta_{ij}|\leq C_{3}<\infty\right\},$
    where $C_{1}$, $C_{2}$, and $C_{3}$ are positive constants.
@@ 4.1 Embedding Error Caused by Misspecified Curvature
Motivated by 2.1, which states that misspecifying curvature leads to the non-existence of distance-preserved mapping, we are interested in characterizing the error bound of the distance matrix for $n$ discrete nodes. Here we focus on the two-dimensional case and formulate the problem as follows: Given (1) two hyperbolic spaces $\mathbb{H}_{2}^{-K}$ and $\mathbb{H}_{2}^{-K^{\prime}}$ with $K\neq K^{\prime}$, and (2) a set of discrete points $\{z_{i}\}_{i=1}^{n}\subset\mathbb{H}_{2}^{-K}$, what is the lower error bound of the following approximation in Equation Equation 11?
$\min_{\{z_{i}^{\prime}\}_{i=1}^{n}\subset\mathbb{H}_{2}^{-K^{\prime}}}||\Theta-\Theta^{\prime}||_{F}^{2}=\min_{\{z_{i}^{\prime}\}_{i=1}^{n}\subset\mathbb{H}_{2}^{-K^{\prime}}}\sum_{i\neq j}\left[d_{\mathbb{H}_{2}^{-K}}({z_{i},z_{j}})-d_{\mathbb{H}_{2}^{-K^{\prime}}}(z_{i}^{\prime},z_{i}^{\prime})\right]^{2}$ (11)
Note that Equation Equation 11 considers the error of pairwise distance matrix, whose value is invariant to the choice of parameterization of hyperbolic geometry. Thus, we can replace the distance metric in Equation 11 to be the metric defined in Equation 2 under the Poincaré disk model on the complex plane, $\mathbb{\widetilde{D}}_{2}$. Utilizing the Schwarz-Pick Theorem (Osserman,, 1999) in complex analysis, we can establish the following lower bound:
@@
Note that a holomorphic function in complex analysis is defined as a complex differentiable mapping such that it can be locally approximated by Taylor expansion. In the context of model optimization where $\Theta$ denotes the true parameters and $\Theta^{\prime}$ denotes the estimators, the holomorphic assumption essentially means that the likelihood function is changing smoothly within a small neighborhood of the true latent positions.
@@
In Chami et al., (2019), the authors proposed the Hyperbolic Graph Convolutional Network (HGCN) model that learns hyperbolic embeddings, finding that HGCN achieved much better performance on downstream tasks with learnable curvature parameters other than fixing them as $-1$. While the authors illustrated such improvement from a numerical perspective, our Theorem 4.1 provides a theoretical justification. Furthermore, while Theorem 4.1 in Chami et al., (2019) states that for distance-based linear classifiers there always exist equivalent sets of embeddings from hyperbolic spaces with different curvatures, our Theorem 4.1 further indicates that the best embeddings can only be obtained when the curvature is optimal.
@@ 4.2 Consistency Results
In this section, we present the consistency results of the maximum-likelihood estimators. Developing consistency results for the maximum-likelihood estimators with hyperbolic geometry is a challenging problem due to the non-linearity and non-convexity of the hyperbolic distance metric. We establish consistency results for two situations: (1) the consistency rates of $\hat{P}$ and $\hat{\Theta}$, where the latent space dimension is allowed to be misspecified; and (2) the consistency rates of $\hat{K}$ and $\hat{Z}$, where the latent space dimension is correctly specified.
We first characterize the consistency of $\hat{P}$ and $\hat{\Theta}$ with potentially misspecified latent space dimension following the approximate low-rank framework in the matrix completion literature (Davenport et al.,, 2014). In our case, the approximate low-rank framework essentially considers matrices with a bounded nuclear norm: $\|Z\Lambda Z^{T}\|_{*}\leq n\sqrt{(d+1)}\cosh(\sqrt{C_{1}}C_{3})$. Here, $*\|\cdot\|_{*}$ represents the summation of the singular values of a matrix.
Theorem 4.2 (Consistency of $\Theta$ and $P$).
Suppose Assumptions I-III hold, and further $\|\hat{Z}\Lambda\hat{Z}^{T}\|_{*}\leq n\sqrt{(d+1)}\cosh(\sqrt{C_{1}}C_{3})$ holds, then regardless of the column dimensions of $Z$ and $\hat{Z}$, we have with probability at least $1-\tilde{C}_{1}/n$,
$\frac{1}{n(n-1)}\|\hat{P}-P\|_{F}^{2}$ $\leq\tilde{C}\cosh{\sqrt{C}_{1}C_{3}}\alpha_{\sigma}\sqrt{\frac{d+1}{n}},$
@@
Theorem 4.2 indicates that both $\|\hat{P}-P\|_{F}^{2}/[n(n-1)]$ and $\|\hat{\Theta}-\Theta\|_{F}^{2}/[n(n-1)]$ have consistency rates of $O_{p}(1/\sqrt{n})$, which hold even if the latent space dimension of the working model is misspecified. Note that when the latent space dimension is correctly specified, the assumption of $\|\hat{Z}\Lambda\hat{Z}^{T}\|_{*}\leq n\sqrt{(d+1)}\cosh(\sqrt{C_{1}}C_{3})$ can be directly induced from Assumption I, since $\|\hat{Z}\Lambda\hat{Z}^{T}\|_{*}\leq\sqrt{(d+1)}\|\hat{Z}\Lambda\hat{Z}^{T}\|_{F}\leq n\sqrt{(d+1)}\cosh(\sqrt{C_{1}}C_{3})$.
Theorem 4.2 holds for any latent space dimension $d\geq 2$. Combining Theorem 4.1 and Theorem 4.2, we are then able to characterize the consistency rates of $K$ and $Z$ when the dimension of the latent space is $d=2$. To do so, we would first characterize the consistency rate of $\hat{Z}\Lambda\hat{Z}^{T}$ and then derive the rate for $\hat{Z}$ with additional regularity assumptions on the singular values of $Z$.
Theorem 4.3 (Consistency of $K$ and $Z$).
Given $d=2$, suppose Assumptions I-III and the assumptions stated in Theorem 4.1 hold, then we have:
$|\hat{K}-K|^{2}=O_{p}(1/\sqrt{n}),\quad\|\hat{Z}-Z\|_{F}^{2}/n=O_{p}(1/\sqrt{n}).$
@@
In Shalizi and Asta, (2017), the authors considered a related problem regarding the consistency results of maximum-likelihood based estimators while making a strong asymptotic assumption on the link probability, which does not fit with our model setup. Furthermore, our results have the following advantages: (1) in Shalizi and Asta, (2017), the consistency results of latent positions are derived assuming the latent space curvature is known, while we consider a more practical situation where both the latent positions $Z$ and the curvature parameter $K$ need to be estimated; and (2) the analytic approach used in Shalizi and Asta, (2017) is not able to characterize the consistency rates of estimators, while in Theorem 4.2 and Theorem 4.3 we derive the consistency rates with techniques in low-rank matrix completion literature.
@@
The consistency rates derived in Theorem 4.3 can be naturally generalized to higher dimensions if Theorem 4.1 holds for any $d\geq 2$. Examining the statement of Theorem 4.1 for higher-dimensional hyperbolic spaces is itself an interesting but challenging mathematical problem in the research of hyperbolic geometry and is left for further exploration.
@@ 5 Simulation Studies
In this section, we showcase simulation results to assess the properties and performance of the proposed model. We focus on the following aspects: (1) the effect of latent space curvature on the properties of generated networks; (2) the estimation consistency of the maximum-likelihood estimators; (3) the impact of latent space curvature on model fitting; and (4) power analysis and coverage rate analysis of a proposed inference procedure on the curvature parameter.
We employ a general setup where networks are generated based on the model defined in (4) with a $2$-dimensional hyperbolic space. The link function is a logistic-type function $\sigma(x)=2\text{logistic}(-x)=2/(1+e^{x})$, such that the edge connecting probability is $1$ when the distance is 00. Latent positions are generated from $i.i.d.$ samples following uniform distribution on the hyperbolic disk with radius $3$, such that the edge connecting probability has a lower bound of around $0.005$. More details of the uniform distribution on the hyperbolic disk as well as the derivation of the sampling procedure can be found in the supplemental material.
@@ 5.1 Model Properties versus Latent Space Curvature
We illustrate the geometric effect of the latent space curvature on model properties with simulated networks, each with $500$ nodes with different choices of $K\in\{0,0.1,1,2,5,10\}$. Here $K=0$ corresponds to the Euclidean latent space, with the same link function and the same distribution of latent positions (i.e., the uniform distribution on the Euclidean disk). Table 1 summarizes the average descriptive statistics with $500$ replicates for each $K$. As the latent space becomes more hyperbolic, i.e., $K$ becomes larger, the simulated networks become more similar to trees that have hierarchical structures. Particularly, when $K$ increases, the edge density and transitivity decrease. At the same time, betweenness centrality, average path length, and diameter increase.
K Edge Density Transitivity Betweenness Centrality Average Path Length Diameter
0 0.208 0.309 0.002 1.793 3.000
0.1 0.194 0.297 0.003 1.809 3.000
1 0.126 0.235 0.006 1.903 3.000
2 0.086 0.181 0.011 2.020 3.000
5 0.040 0.091 0.036 2.409 4.000
10 0.021 0.039 0.066 2.855 5.058
Table 1: Average descriptive statistics of simulated networks with different curvatures.
We also examine the node degree distributions of four simulated networks, each with $5000$ nodes and $K\in\{0,0.1,1,10\}$. As illustrated in Figure 6, when $K$ becomes larger, the node degree distribution becomes skewer and more similar to the power-law distribution commonly observed in real-world social and internet networks. From both experiments above, we can see that the curvature parameter has a great effect on the global and local properties of generated networks, where a greater curvature parameter leads to more heterogeneous network structures.
Figure 6: Node degree distribution of simulated networks with different curvatures.
@@ 5.3 Embedding Error versus Latent Space Curvature
In the second experiment, we generate networks with $1000$ nodes from the hyperbolic model while using Euclidean models to fit them, with increasing latent space dimensions from $2$ to $20$. We select curvature parameter values for two situations: $K=0.1$, where the latent space is closer to Euclidean (where $K=0$); and $K=1$, a significantly hyperbolic latent space. Figure 10 reports the relative error of the estimated probability matrix of different models, averaged over $100$ replicates. In both settings, the hyperbolic model achieves the lowest estimation error since it has the learnable curvature parameter. When $K=0.1$, the performance of $E2$ is comparable to that of $H2$, whereas Euclidean models with higher dimensions show larger estimation errors attributed to overfitting. In the case of $K=1$, $E2$ experiences a considerably high estimation error. Moreover, other Euclidean models with greater latent space dimensions exhibit nearly twice the error of the hyperbolic model. These results demonstrate that the influence of latent space curvature on model fitting differs from, and cannot be substituted by, the impact of latent space dimension.
[…]
@@ 5.4 Statistical Inference of the Curvature Parameter $K$
The curvature parameter $K$ characterizes how heterogeneous a network is, which can be reflected from its global properties (Table 1 and Figure 6). An interesting practical question is how to conduct statistical inference over the curvature parameter. We propose a two-step procedure: (1) conduct a likelihood-ratio test to determine whether the latent space geometry is Euclidean ($K=0$) or hyperbolic ($K>0$); and (2) construct a confidence interval for $K$ if the testing result indicates $K>0$.
Likelihood Ratio Test of $K$
To determine whether the latent space geometry is hyperbolic or Euclidean, we propose a likelihood ratio test for the hypothesis testing problem: $H_{0}:~{}K=0\mbox{ v.s. }H_{1}:~{}K>0$, where under $H_{0}$ the latent space geometry is Euclidean and under $H_{1}$ it is hyperbolic. The test statistic is defined as
$\Lambda(A)=-2\log\frac{\max_{K=0}L(K,Z;A)}{\max_{K\geq 0}L(K,Z;A)}=-2\log\frac{\max_{K=0}L(K,Z;A)}{\max\{L(K=0,Z;A),\max_{K>0}L(K,Z;A)\}},$
where $L$ is the model likelihood function. The numerator in the equation above is the likelihood of the Euclidean model, while the denominator is the maximum likelihood of the Euclidean model and the hyperbolic model. By Euclidean model, we refer to the distance latent space model that takes the same form of (4) while replacing the hyperbolic distance metric with the Euclidean one of the same dimension. Given the challenging nature of deriving the asymptotic distribution of the likelihood ratio test statistic in the context of the complex nonlinear latent space model, we employ a parametric bootstrap method to conduct inference in practice.
The parametric bootstrap likelihood ratio test operates as the following. Initially, we compute the test statistic $\hat{\Lambda}(A)$ by fitting both the Euclidean and hyperbolic models using the given network data. Subsequently, we generate bootstrap network samples $A_{1},...,A_{b},...,A_{B}$ under the Euclidean model (the model under $H_{0}$), with the estimated Euclidean latent positions serving as the “true” parameters. For each of the bootstrap network samples $A_{b}$, we calculate the test statistics $\hat{\Lambda}(A_{b})$ by fitting both Euclidean and hyperbolic models. Finally, the $p$-value is reported as the proportion of $\{\hat{\Lambda}(A_{b})\}_{b=1}^{B}$ that are greater than $\hat{\Lambda}(A)$, providing a measure of the statistical significance of the observed test statistic.
We perform an experiment to assess the effectiveness of the proposed testing procedure. The experiments are conducted with varying numbers of nodes $n\in\{100,200,300,400,500\}$, and different values of the curvature parameter $K\in\{0,0.01,0.05,0.1\}$. For $K=0$, the networks are generated under the Euclidean model, representing the null hypothesis model. In contrast, for $K\in\{0.01,0.05,0.1\}$, the networks are generated under the alternative hypothesis hyperbolic model. We set the significance level to be $0.05$ and for each network, $200$ bootstrap samples are simulated. The left plot in Figure 11 presents the proportion of the $100$ experiments where $H_{0}$ is rejected for each experimental setting. We can see that when $K=0$, the Type-I error rates are well controlled. Under the alternative hypotheses, the proposed test achieves full power as $K$ deviates away from 00 with increasing $n$.
Confidence Interval of $K$
Extending the parametric bootstrap idea, we can also construct a confidence interval for $K$ given that $K>0$. The procedure is as follows. Given a network $A$, we fit the hyperbolic latent space model and obtain estimators $\hat{K}$ and $\hat{Z}$. Following this, we simulate bootstrap network samples $A_{1},...,A_{b},...,A_{B}$ with $\hat{K}$ and $\hat{Z}$ serving as the “true” parameters. For each bootstrap network sample $A_{b}$, we obtain the estimated curvature parameter $\hat{K}_{b}$. Finally, a confidence interval can be constructed based on the bootstrap estimators using the empirical distribution of bootstrap samples. To assess the numerical performance of the bootstrap confidence interval, we generate networks with the number of nodes ranging from $n\in\{100,200,300,400,500\}$ and different curvature parameters $K\in\{0.1,0.5,1,1.5\}$. We set the bootstrap sample size to be $200$ and repeat the procedure $100$ times. The right plot of Figure 11 presents the coverage rates of $95\%$ confidence intervals, where the coverage rates closely resemble theoretical expectations, particularly as $n$ increases.
Figure 11: Left: Reject Proportion of the Bootstrap Testing Procedures; Right: Coverage Proportion of the Bootstrap Confidence Intervals.

### v2

@@ 4 Theoretical Analysis
In this section, we first utilize hyperbolic geometry techniques to characterize the error bound for the embedded distance matrix of latent positions from hyperbolic spaces with different curvatures, which reveals the importance of the latent space curvature for general distance-based hyperbolic embedding methods. We then present the consistency rates for the maximum likelihood estimators of both the curvature and the latent positions.
We use $P$, $Z$, and similar symbols, to denote the true values of parameters, and use $\hat{P}$, $\hat{Z}$, and so on, to represent their estimators. The following assumptions are made throughout.
@@
    Values of both the estimators and true model parameters belong to the bounded parameter space $\Xi=\left\{0<C_{1}^{-1}\leq K\leq C_{1},0<C_{2}\leq|\Theta_{ij}|\leq C_{3}<\infty\right\},$ where $C_{1}$, $C_{2}$, and $C_{3}$ are positive constants.
@@ 4.1 Embedding Error Caused by Misspecified Curvature
Motivated by 2.1, which states that misspecifying curvature leads to the non-existence of distance-preserved mapping, we are interested in characterizing the error bound of the distance matrix for $n$ discrete nodes. Here we focus on the two-dimensional case and formulate the problem as follows. Given (1) two hyperbolic spaces $\mathbb{H}_{2}^{-K}$ and $\mathbb{H}_{2}^{-K^{\prime}}$ with $K\neq K^{\prime}$, and (2) a set of discrete points $\{z_{i}\}_{i=1}^{n}\subset\mathbb{H}_{2}^{-K}$, what is the lower error bound of the following approximation in Equation 8?
[…]
Note that Equation 8 considers the error of pairwise distance matrix, whose value is invariant to the choice of parameterization of hyperbolic geometry. Thus, we can replace the distance metric in Equation 8 to be the metric defined in Equation 2 under the Poincaré disk model on the complex plane, $\mathbb{\widetilde{D}}_{2}$. Utilizing the Schwarz-Pick Theorem (Osserman, 1999) in complex analysis, we can establish the following lower bound:
@@
Note that a holomorphic function in complex analysis is defined as a complex differentiable mapping such that it can be locally approximated by Taylor expansion. Such smoothness assumption aligns with the nature of the minimization problem in Equation 8, where the loss function itself is inherently smooth. In the context of model optimization where $\Theta$ denotes the true parameters and $\Theta^{\prime}$ denotes the estimators, the holomorphic assumption essentially means that the likelihood function is changing smoothly within a small neighborhood of the true latent positions.
@@
In Chami et al. (2019), the authors proposed the Hyperbolic Graph Convolutional Network (HGCN) model that learns hyperbolic embeddings, finding that HGCN achieved much better performance on downstream tasks with learnable curvature parameters other than fixing them as 1-1. While the authors illustrated such improvement from a numerical perspective, our Theorem 4.1 provides a theoretical justification. Furthermore, while Theorem 4.1 in Chami et al. (2019) states that for distance-based linear classifiers there always exist equivalent sets of embeddings from hyperbolic spaces with different curvatures, our Theorem 4.1 further indicates that the best embeddings can only be obtained when the curvature is optimal.
@@ 4.2 Consistency Results
In this section, we present the consistency results of the maximum-likelihood estimators. Developing consistency results for the maximum-likelihood estimators with hyperbolic geometry is a challenging problem due to the non-linearity and non-convexity of the hyperbolic distance metric. We establish consistency results for two situations: (1) the consistency rates of $\hat{P}$ and $\hat{\Theta}$, and (2) the consistency rates of $\hat{K}$ and $\hat{Z}$, where the latent space dimension is correctly specified.
We first characterize the consistency of $\hat{P}$ and $\hat{\Theta}$ following the approximate low-rank framework in the matrix completion literature (Davenport et al., 2014). In our case, we refer to the matrices as having a bounded nuclear norm: $\|Z\Lambda Z^{T}\|_{*}\leq n\sqrt{(d+1)}\cosh(\sqrt{C_{1}}C_{3})$. Here, $\|\cdot\|_{*}$ represents the summation of the singular values of a matrix.
Theorem 4.2 (Consistency of $\hat{\Theta}$ and $\hat{P}$).
Suppose Assumptions I-III hold, then we have with probability at least $1-\tilde{C}_{1}/n$,
$\frac{1}{n(n-1)}\|\hat{P}-P\|_{F}^{2}$ $\leq\tilde{C}\cosh(\sqrt{C}_{1}C_{3})\alpha_{\sigma}\sqrt{\frac{d+1}{n}},$
@@
Theorem 4.2 indicates that both $\|\hat{P}-P\|_{F}^{2}/[n(n-1)]$ and $\|\hat{\Theta}-\Theta\|_{F}^{2}/[n(n-1)]$ have consistency rates of $O_{p}(1/\sqrt{n})$. Note that the approximate low-rank condition can be induced from Assumption I, since $\|\hat{Z}\Lambda\hat{Z}^{T}\|_{*}\leq\sqrt{(d+1)}\|\hat{Z}\Lambda\hat{Z}^{T}\|_{F}\leq n\sqrt{(d+1)}\cosh(\sqrt{C_{1}}C_{3})$.
Theorem 4.2 holds for any latent space dimension $d\geq 2$. Combining Theorem 4.1 and Theorem 4.2, we are then able to characterize the consistency rates of $\hat{K}$ and $\hat{Z}$ when the dimension of the latent space is $d=2$. To do so, we would first characterize the consistency rate of $\hat{Z}\Lambda\hat{Z}^{T}$ and then derive the rate for $\hat{Z}$ with additional regularity assumptions on the singular values of $Z$.
Theorem 4.3 (Consistency of $\hat{K}$ and $\hat{Z}$).
Given $d=2$, suppose Assumptions I-III and the assumptions stated in Theorem 4.1 hold, then we have $|\hat{K}-K|^{2}=O_{p}(1/\sqrt{n})$ and $\|\hat{Z}-Z\|_{F}^{2}/n=O_{p}(1/\sqrt{n}).$
@@
In Shalizi and Asta (2017), the authors considered a related problem regarding the consistency results of maximum-likelihood based estimators while making a strong asymptotic assumption on the link probability, which does not fit with our model setup. Furthermore, our results have the following advantages: (1) in Shalizi and Asta (2017), the consistency results of latent positions are derived assuming the latent space curvature is known, while we consider a more practical situation where both the latent positions $Z$ and the curvature parameter $K$ need to be estimated; and (2) the analytic approach used in Shalizi and Asta (2017) is not able to characterize the consistency rates of estimators, while in Theorem 4.2 and Theorem 4.3 we derive the consistency rates with techniques in low-rank matrix completion literature.
@@
The consistency rates derived in Theorem 4.3 can be naturally generalized to higher dimensions if Theorem 4.1 holds for any $d\geq 2$. Examining the statement of Theorem 4.1 for higher-dimensional hyperbolic spaces is itself an interesting but challenging mathematical problem in hyperbolic geometry and is left for further exploration.
@@ 5 Simulation Studies
In this section, we use simulation studies to assess the properties and performance of the proposed model. We focus on the following aspects: (1) the effect of latent space curvature on the properties of generated networks; (2) the estimation consistency of the maximum-likelihood estimators; (3) the impact of latent space curvature on model fitting; and (4) power analysis and coverage rate analysis of a proposed inference procedure on the curvature parameter. We employ a general setup where networks are generated based on the model defined in (4) with a $2$-dimensional hyperbolic space. The link function is a logistic-type function $\sigma(x)=2\text{logistic}(-x)=2/(1+e^{x})$, such that the edge connecting probability is $1$ when the distance is 0. Latent positions are generated from $i.i.d.$ samples following uniform distribution on the hyperbolic disk with radius $3$, such that the edge connecting probability has a lower bound of around $0.005$. More details of the uniform distribution on the hyperbolic disk as well as the derivation of the sampling procedure can be found in the Supplementary Materials.
@@ 5.1 Model Properties versus Latent Space Curvature
We first illustrate the geometric effect of the latent space curvature on the generated networks. From both experiments presented in the Supplementary Materials (Section B.1), we can see that the curvature parameter has a great effect on the global and local properties of generated networks. Specifically, a larger $K$ leads to lower edge density and more heterogeneous network structures. Therefore, in practice, $K$ can be interpreted as a modeling parameter for sparsity and heterogeneity, where larger values of $K$ indicate sparser and more heterogeneous networks.
@@ 5.3 Embedding Error versus Latent Space Curvature
In the second experiment, we generate networks with $1000$ nodes from the hyperbolic model while using Euclidean models to fit them, with increasing latent space dimensions from $2$ to $20$. We conduct experiments with increasing curvature parameters, from $K=0.1$, where the latent space is closer to Euclidean, to $K=3$, where the latent space is significantly hyperbolic. Figure 8 reports the relative error of the estimated probability matrix of different models, averaged over $100$ replicates. In all settings, the hyperbolic model achieves the lowest estimation error since it has the learnable curvature parameter. When $K=0.1$, the performance of $E2$ is comparable to that of $H2$, whereas Euclidean models with higher dimensions show larger estimation errors attributed to overfitting. In the cases of $K=1,2$, and $3$, $E2$ experiences a considerably high estimation error. Moreover, other Euclidean models with greater latent space dimensions exhibit nearly twice the error of the hyperbolic model. These results demonstrate that the influence of latent space curvature on model fitting differs from, and cannot be substituted by, the impact of latent space dimension, and these two effects are not interchangeable (see Section 5.5).
Figure 8: Relative error of the probability matrix, estimated by the hyperbolic model and Euclidean models with different latent space dimensions. “H2” represents the two-dimensional hyperbolic model and “Ex{x}” represents the x{x}-dimensional Euclidean model.
@@ 5.4 Statistical Inference of the Curvature Parameter $K$
The curvature parameter $K$ characterizes how heterogeneous a network is, which can be reflected from its global properties as shown in Section 5.1. An interesting practical question is how to conduct statistical inference over the curvature parameter. We propose a two-step procedure: (1) conduct a likelihood-ratio test to determine whether the latent space geometry is Euclidean ($K=0$) or hyperbolic ($K>0$); and (2) construct a confidence interval for $K$ if the testing result indicates $K>0$. Given the challenging nature of deriving the asymptotic distribution of the likelihood ratio test statistic in the context of the complex nonlinear latent space model, we employ a parametric bootstrap method to conduct inference in practice. Due to space limitations, the procedure and simulation results are presented in the Supplementary Materials.
5.5 Impact of Latent Space Dimension on Model Estimation
We also investigate the impact of mis-specifying the latent space dimension on model estimation through simulation studies, finding that (1) under-specifying latent space dimension is more detrimental to model fitting than over-specifying it; and (2) under-specifying latent space dimension would lead to over-estimating $K$, and vice versa. These results further enhance our understanding of the interplay between latent space curvature and dimension. Detailed results are provided in the Supplementary Materials (Section B.4).